OpenStudy (vera_ewing):

Graph the first six terms of a sequence where a1 = -10 and d = 3.

OpenStudy (vera_ewing):

A.

OpenStudy (vera_ewing):

B.

OpenStudy (vera_ewing):

C.

OpenStudy (vera_ewing):

D.

OpenStudy (misssunshinexxoxo):

@jim_thompson5910

OpenStudy (zzr0ck3r):

what is d?

OpenStudy (zzr0ck3r):

no no. I dont understand the question. But I don't know much about arithmetic sequences. Is that what you are doing?

OpenStudy (freckles):

is it arithmetic sequence though ? I assume it is because a lot of people use d to mean common difference but I'm pretty sure it is not universal notation.

OpenStudy (vera_ewing):

Yes it's an arithmetic sequence.

OpenStudy (freckles):

if it is an arithmetic sequence, the sequence goes like this: \[a_1,a_1+d,a_1+2d,a_1+3d,...\]

OpenStudy (freckles):

a1 is the first term a1+d is the second term ... a1+(n-1)d is the nth term

OpenStudy (freckles):

ok so a1=-10 and you know a2=?

OpenStudy (freckles):

well a1+d isn't -20 Ok I'm going to use the x-axis as the n-axis and the y-axis as the outputs

OpenStudy (freckles):

a1+d=-10+3=-7

OpenStudy (freckles):

so graphing (1,-10) and (2,-7)

OpenStudy (vera_ewing):

Oh so it's A.

OpenStudy (freckles):

 I ran out of room so I pretending -10 is someone down there